Aluminum Carbonate has a density of 1.770 g/cm^3. Calculate the mass (in pg) of oxygen atoms present in the cubic crystal of AlCar that has a volume of 457.6 Å^3
1
Expert's answer
2020-10-14T05:48:29-0400

Volume is 457 .6 * 10^-24 cm^3 .


density = mass / volume .


so , mass = volume * density . = 457.6 * 10^-24 * 1.770 = 8.09 * 10^-22 g .


So , Mass = 8.09 * 10^-10 pg . Answer .